# TextMine: LLM-Powered Knowledge Extraction for Humanitarian Mine Action

### v1

@@ Abstract
Humanitarian Mine Action has generated extensive best-practice knowledge, but much remains locked in unstructured reports. We introduce TextMine, an ontology-guided pipeline that uses Large Language Models to extract knowledge triples from HMA texts. TextMine integrates document chunking, domain-aware prompting, triple extraction, and both reference-based and LLM-as-a-Judge evaluation. We also create the first HMA ontology and a curated dataset of real-world demining reports. Experiments show ontology-aligned prompts boost extraction accuracy by 44.2%, cut hallucinations by 22.5%, and improve format conformance by 20.9% over baselines. While validated on Cambodian reports, TextMine can adapt to global demining efforts or other domains, transforming unstructured data into structured knowledge.
TextMine: LLM-Powered Knowledge Extraction for Humanitarian Mine Action
@@ 1 Introduction
Landmine clearance remains a critical humanitarian challenge. In 2022 there were 4,710 casualties globally, 85% civilians (United Nations, 2025; Inclusion, 2023). Mine action authorities worldwide publish technical reports, surveys, lessons learned. However, this demining knowledge is largely underutilized due to unstructured formats and the absence of an integrated demining knowledge base. Extracting structured knowledge in the form of a knowledge graph from these reports allows sharing, accessing, and learning from the recorded experiences. It will provide actionable insights to improve decision-making and operational efficiency in humanitarian demining.
Cambodia has suffered over 65,000 casualties since 1979 (United Nations Development Programme (2021), UNDP). In collaboration with the Cambodian Mine Action Centre (CMAC), we aim to convert their technical reports into a structured knowledge base. An example context and corresponding extracted triples under our HMA ontology is provided in Figure 1.
@@
Figure 1: Example of extracting KG triples from humanitarian mine action (HMA) texts, guided by specific ontology. The task is to extract as many triples as possible while ensuring conformance to the ontology and faithfulness to the source text.
Constructing knowledge graphs from unstructured text remains resource–intensive. Traditional pipelines, such as Named Entity Recognition followed by Relation Extraction, require separate models and often domain‐specific fine‐tuning (Nadeau and Sekine, 2009; Pawar et al., 2017). Task‐independent LLMs can automate both subtasks without additional training and match or exceed fine‐tuned models (Wang et al., 2023; Jinensibieke et al., 2024). Recent benchmarks like Text2KGBench (Mihindukulasooriya et al., 2023) and ontology‐guided sentence‐level methods (Cauter and Yakovets, 2024) have demonstrated promise, but they rely on toy ontologies and single‐sentence inputs, limiting their applicability to complex, domain‐specific documents.
Unlike prior sentence level approaches, TextMine reasons over entire paragraphs, enabling coreference resolution and multi step inference, while relying on a practical operational HMA ontology that is orders of magnitude larger than those used in existing benchmarks (Table 1).
Table 1: Comparison of TextMine with other triple extraction methods. “Operational Ontology” denotes active use in real world applications.
Criteria Text2KGBench (Cauter and Yakovets, 2024) TextMine
Domain General Maintenance HMA
Input Granularity Sentence Phrase Paragraph
Operational Ontology
Paragraph Level
One key challenge in our work is to investigate effective methods to activate LLMs to reason over paragraph-level text while adhering to a given ontology. Another major challenge is to establish a reliable evaluation framework for the extracted triples. We summarize the contributions of this work as follows:
  • Automated Demining Knowledge Extraction Framework. We introduce TextMine, a prompt-based pipeline that combines layout-aware document chunking, ontology-guided extraction, and multi-perspective evaluation. To our knowledge, this is the first LLM application of knowledge extraction in the HMA domain.
  • HMA Ontology and Evaluation Dataset. We introduce a dedicated ontology for humanitarian demining operations, systematically categorizing operational entities and relationships. Alongside this ontology, we create a human-annotated, ontology-aligned triple dataset, filling a critical resource gap for this domain.
  • Reference-Based and Reference-Free Evaluation. We evaluate extracted triples against our annotated dataset and introduce a bias-aware LLM-as-Judge framework for reference-free scoring. Experiments on closed and open LLMs show that position bias skews rankings, which randomizing output order mitigates for some models.
  • In-Context Learning Optimization. We examine demonstration design and find that prompts enriched with ontology-aligned examples improve triple extraction accuracy by up to 44.2%, reduce hallucinations by 22.5%, and enhance format conformance by up to 20.9% compared to baseline prompts. These findings provide practical insights for prompt construction.
2 The Humanitarian Mine Action
Figure 2: Left: CMAC landmine clearance operation. Right: CMAC Technical Progress reports. Our workflow extracts knowledge from demining reports with the aim to leverages it to guide future clearance planning.
Demining operations in Cambodia incur costs in the billions (Harris, 2000), yet most agencies still manage data via spreadsheets and manual review, which impedes timely analysis and collaboration. In partnership with CMAC, we address this by automating knowledge extraction from both field and technical reports (Figure 2). Our framework serves as a proof of concept for LLM driven demining knowledge extraction, transforming unstructured reports into structured insights that can directly inform and optimize future clearance planning, and is readily adaptable to other mine affected regions.
2.1 Task Description
Knowledge triple extraction from HMA documents can be formally defined as follows: Given an ontology $\mathcal{O}=(\mathcal{E},\mathcal{R})$, where $\mathcal{E}$ represents ontology entities and $\mathcal{R}$ represents ontology relations, and a textual context $C$, the objective is to design a prompt $P(C,\mathcal{O})$ that guides an extractor model $M$ to extract triples $T=\{(s_{1},r_{1},o_{1}),(s_{2},r_{2},o_{2}),\cdots,(s_{n},r_{n},o_{n})\mid s,o\in\mathcal{E},\,r\in\mathcal{R}\}$, where $n$ is the number of extracted triples ($n=|T|$), subjects and objects from source text $s,o\in\mathcal{E}$ (mapped to the entities $\mathcal{E}$ in the ontology) and relations from the ontology $r\in\mathcal{R}$. The extracted triples must remain consistent with both the ontology and the source text. Shortly, this problem can be written as $T=M\bigl(P(C,\mathcal{O})\bigr)$.
@@ 3 TextMine Overview
Figure 3: TextMine Overview. Reports are preprocessed into paragraph chunks, used as test inputs during inference. Each chunk is combined with an instruction template and ontology, then passed to LLMs for triple extraction. We apply a multi-perspective evaluation using both reference-based and reference-free methods. Extracted triples can be stored in a Graph DB for querying by developers and domain experts.
@@
3.1 Layout-Aware Document Chunking
@@
3.2 Ontology-Guided Knowledge Extraction
Given the text chunks as input, our goal is to extract triples $T=\{(s_{1},r_{1},o_{1}),\cdots,(s_{n},r_{n},o_{n})\mid s,o\in\mathcal{E},\,r\in\mathcal{R}\}$, where $s,o\in\mathcal{E}$ (subject and object pairs) and $r\in\mathcal{R}$ (relations). First, the extracted triples must be accurate w.r.t., the source text. An ideal extraction system would extract all triples (recall=1) precisely (precision=1). Second, extracted triples must conform to a specified ontology so that extracted demining knowledge can be stored and shared between organizations and countries in a compatible way. We address this by combining a domain-specific HMA ontology with LLM in-context learning to extract triples from paragraphs.
HMA Ontology
We performed a survey of demining related ontologies together with domain experts and we incorporate six ontologies from IMSMA [^1] Information Management System for Mine Action (IMSMA), https://www.gichd.org/our-response/information-management/imsma-core/.
These ontologies are used for demining information system but not strictly for HMA, We then manually filter out concepts and relations unrelated to HMA. Furthermore, we add a more general humanitarian domain ontology from Empathi (Gaur et al., 2019) to make the overall HMA ontology more comprehensive. As a result, HMA ontology integrates seven ontologies (160 entity, 86 relation types) covering diverse aspects of humanitarian demining action (see Table 2).
Table 2: Number of entity types and relation types for each ontology.
Empathi Assistance Accident Activity Land Location Victim
Entity Types 20 14 34 29 31 5 27
Relation Types 22 7 6 24 10 5 12
@@ 3.2 Ontology-Guided Knowledge Extraction
We design five prompting strategies for knowledge triple extraction: (1) Zero-shot (instruction and context only), (2) One-shot with Random Sentence (RS), (3) One-shot with Random Paragraph (RP), (4) One-shot with Ontology-Aligned Sentence (OS), and (5) One-shot with Ontology-Aligned Paragraph (OP). One example prompt is provided in Appendix TextMine: LLM-Powered Knowledge Extraction for Humanitarian Mine Action.
Ontology-aligned demonstrations share the same ontology (entity types and relationships) as the target context, while random demonstrations use unrelated ontology. Sentences are extracted from paragraphs using NLTK’s sentence tokenizer [^2] https://www.nltk.org/howto/stem.html. For OS/OP prompts, we design a retrieval algorithm that selects demonstrations from our annotated dataset (Section 4.1) by identifying the shortest context-answer pair matching the target ontology. This is for minimizing token costs while retaining high semantic similarity. To prevent data leakage and ensure a fair evaluation, we implement a second retrieval step if the initially retrieved demonstration contains the same or part of target context as the test instance. In such cases, we select the next shortest matching example instead. This ensures that retrieved demonstrations do not overlap with the evaluation context, preserving the integrity of the inference.
These prompt templates are designed to test our two hypotheses: (a) Ontology alignment enhances accuracy by priming the model with ontology-specific reasoning. Semantically aligned demonstrations help constrain the label space and improve precision (Min et al., 2022; Long et al., ). This effect is evident in our results as the contrast between RS/RP and OS/OP confirms the benefit of ontology alignment. (b) Paragraph-level context improves extraction performance by providing richer demonstrations that reflect how entities and relations are introduced across sentences in real-world reports. However, our results do not support this hypothesis as comparisons between RS vs. RP and OS vs. OP show no consistent improvement from paragraph-level context.
@@ 3.3 Multi-Perspective Evaluation
Humanitarian Demining is a high-stakes decision-making domain, incorrect triples, especially hallucinated ones, can misinform demining operations, leading to safety risks. A comprehensive evaluation of triple extraction requires assessing accuracy, reliability, and structural validity. We evaluate models across three dimensions: (1) Triple Extraction Accuracy, (2) Hallucination Rate, and (3) Format Conformance.
@@
We employ N-gram matching-based metrics such as BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), and METEOR (Banerjee and Lavie, 2005) to assess the extracted triples. Additionally, we incorporate BERTScore (Zhang et al., 2019), which leverages word embeddings to capture semantic similarity beyond lexical overlap. To enhance the accuracy of our metrics, we first apply stemming and lemmatization using NLTK to normalize morphological variations. We then compute the accuracy metrics against the manually annotated test set.
Triple extraction accuracy serves as the primary metric, as it directly measures how well models extract knowledge triples from text. However, accuracy alone does not fully capture model reliability. The hallucination rate evaluates faithfulness by detecting extraneous or fabricated information, while format conformance ensures that outputs adhere to a syntactically valid structure, enabling seamless integration into downstream applications.
@@
Accuracy primarily measures how well the extracted triples match reference triples, but it does not fully assess whether the generated content is grounded in the input. A model could produce plausible triples that are semantically similar to the original text (thus achieving a high accuracy score) yet still be incorrect, meaning they do not appear in the input but seem reasonable. This distinction is crucial in high-stakes applications like humanitarian demining, where misleading information can have serious consequences. Hallucination is a prevalent issue in LLM-generated content (Ji et al., 2023; Huang et al., 2023; Xu et al., 2024) and a critical aspect of our evaluation. To quantify this, for each extracted triple $t=(s,r,o)$ we first normalize $s$, $r$, and $o$, as well as the entire input context (tokenization, lemmatization, lowercasing, and punctuation removal). We then check whether the normalized subject $s$ and object $o$ occur as contiguous substrings in the normalized report text—if not, $t$ is flagged as a hallucination. Similarly, if the normalized relation $r$ is not found in the normalized ontology relation set, $t$ is also flagged. This procedure ensures that even “plausible” but unsupported triples are detected and penalized, thereby maintaining faithfulness and trustworthiness in our system.
@@
Format conformance metric assesses whether the generated triples adhere to the correct syntactic format of $r(s,o)$, where $r$ represents the relation and $s$ and $o$ denote the subject and object, respectively. We consider a triple well-formatted if it follows this structure. We accommodate edge cases where the subject or object contains numerical values with commas, such as hasReliabilityInfo(2,500,011 square meters, landmine/ERW affected areas), or phrases in parentheses, such as hasAccidentOrganisationInfo(Quality of Life Survey (QLS), Department of Victim Assistance of CMAA). Format conformance ensures that extracted triples follow a structured format necessary for practical use. A model with high accuracy but poor format conformance may fail to produce usable outputs, limiting its applicability in real-world knowledge extraction tasks.
@@
where $S^{\prime}$ represents the normalized metric values within $[0,1]$, and the hallucination rate is inverted to penalize higher hallucination. $k=5$ is the number of metrics included in the score. Format conformance is excluded in the Combined Score, as our experimental results show consistently high format conformance across all extraction models and prompt settings, making it non-differentiating. Combined Score provides a holistic measure of extraction quality while mitigating scale differences among individual metrics.
@@
Evaluating generated texts is particularly challenging in domains like humanitarian demining, where annotated datasets are scarce. Demining reports are highly technical and domain-specific, requiring extracted triples to align with predefined ontologies of landmine types, clearance operations, and affected areas. Constructing a manually labeled test set is time-consuming and resource-intensive, limiting large-scale reference-based evaluation. Given these constraints, we explore an LLM-as-a-Judge approach as a potential reference-free evaluation framework for evaluating extracted triples.
LLM-as-a-Judge offers a potential alternative to evaluation when ground-truth data is limited (Friel and Sanyal, 2023; Saad-Falcon et al., 2023; Es et al., 2023). The ultimate objective of our approach is to find a optimal judge LLM setting where the LLM consistently identifies the best candidate answer and provides a reasoned justification for its decision.
We try to find the optimal LLM Judge setting by conducting systematic ranking experiments and analyzing correlations between the LLMs judged ranking and reference-based rankings. For these ranking experiments, we design Judge Prompts that instruct LLMs on evaluation criteria. We use five models including Mistral-7B [^3] https://ollama.com/library/mistral:7b, Llama3-8B [^4] https://ollama.com/library/llama3:8b, Gemma2-9B [^5] https://ollama.com/library/gemma2:9b, LLaMA3-70B [^6] https://ollama.com/library/llama3:70b and GPT-4o [^7] https://openai.com/index/hello-gpt-4o/ as extraction models. We rank five responses from five models using GPT-4o, Llama3.1-70B [^8] https://ollama.com/library/llama3.1:70b, and Llama3.3-70B [^9] https://ollama.com/library/llama3.3:70b as our judge models. [^10] In our experiments, GPT-4o is used via the Azure OpenAI API, all the other open-source models are used via local deployment using Ollama. The judge prompts follow a fixed template with seven placeholders, where the ontology placeholder represents entity and relation types. Formally, let the input set for the LLM judge prompts be $\{O,C,R_{m_{1}},R_{m_{2}},R_{m_{3}},R_{m_{4}},R_{m_{5}}\}$, where $O$ is the ontology set, $C$ is the set of test contexts, and $R_{m_{1}},\cdots,R_{m_{5}}$ are the five sets of candidate answers from the five extractor models. The judge LLM produces a verdict (output as a ranking):
@@
Once the optimal judge LLM setting is determined, we adopt it as the reference-free evaluator to identify the best answer from each extraction, leveraging its reasoning process. Detailed prompt templates and an example of the reasoning process used for evaluation are provided in Appendix TextMine: LLM-Powered Knowledge Extraction for Humanitarian Mine Action.

### v2

# TextMine: Data, Evaluation Framework and Ontology-guided LLM Pipeline for Humanitarian Mine Action

@@ Abstract
Humanitarian Mine Action (HMA) addresses the challenge of detecting and removing landmines from conflict regions. Much of the life-saving operational knowledge produced by HMA agencies is buried in unstructured reports, limiting the transferability of information between agencies. To address this issue, we propose TextMine: the first dataset, evaluation framework and ontology-guided large language model (LLM) pipeline for knowledge extraction in the HMA domain. TextMine structures HMA reports into (subject, relation, object)-triples, thus creating domain-specific knowledge. To ensure real-world relevance, we created the dataset in collaboration with Cambodian Mine Action Center (CMAC). We further introduce a bias-aware evaluation framework that combines human-annotated triples with an LLM-as-Judge protocol to mitigate position bias in reference-free scoring. Our experiments show that ontology-aligned prompts improve extraction accuracy by up to 44.2%, reduce hallucinations by 22.5%, and enhance format adherence by 20.9% compared to baseline models. We publicly release the dataset and code. [^1] https://anonymous.4open.science/r/TextMine_Code_Dataset-D974
TextMine: Data, Evaluation Framework and Ontology-guided LLM Pipeline for Humanitarian Mine Action
@@ 1 Introduction
Humanitarian Mine Action (HMA)—detecting and removing landmines from past (and ongoing) conflicts in order to return land to civilian use—remains a critical humanitarian challenge: in 2022 alone, there were 4,710 casualties globally, 85% of which were civilians (United Nations, 2025; Inclusion, 2023). Over the decades, HMA authorities have published large amount of life-saving knowledge, yet much of it remains locked away in unstructured, free-form text reports, thus making this knowledge largely inaccessible. The automatic extraction and organization of this important information is, therefore, not only a technical advancement, but also a humanitarian imperative that can make life-saving insights more accessible, actionable and transferable across many humanitarian agencies.
@@
Figure 1: Example knowledge graph triple extraction from texts, guided by the HMA ontology. The task is to extract as many triples as possible while ensuring ontology conformance and source-text faithfulness.
To make HMA knowledge accessible, we turn to Information Extraction (IE). Prior research has shown that structuring information from natural language text data is useful for many downstream tasks, such as question answering Xiong et al. (2024), fact retrieval Han et al. (2023) or predicting new facts that were not present originally in the data Broscheit et al. (2020). More importantly, IE pipelines are useful frameworks for quick information exchange between organizations Poletto et al. (2021), which is of vital importance for HMA, because the work is typically done in various agencies from different countries.
To effectively structure such domain-specific information, researchers have proposed data, benchmarks and methods for IE in various domains, including finance Hamad et al. (2024), material science Cheung et al. (2024), medicine Romero et al. (2025) and humanitarian crisis response Fekih et al. (2022). Although LLMs have become ubiquitous, it has been shown that using off-the-shelf LLMs for domain-specific IE is not an optimal approach Pang et al. (2024); Farzi et al. (2024); Dagdelen et al. (2024). Therefore, there is a need for domain-specific IE applications.
In this work, we propose TextMine: the first dataset, evaluation framework and pipeline for IE for the HMA domain. The task is to extract information from HMA reports in the form of (subject, relation, object)-triples, thus creating structured knowledge for the HMA domain (Fig. 1). Prior data and benchmarks (Mihindukulasooriya et al., 2023, inter alia) rely on toy data schemas (i.e., ontologies), which limits their applicability to the complex and specialized HMA documents. By contrast, we created the data in collaboration with the Cambodian Mine Action Center (CMAC). CMAC has been initiated by the UN more than 30 years ago, is one of the world’s most experienced and effective demining organizations. Its collaborations within the UN and the Geneva International Centre for Humanitarian Demining (GICHD), ensure that the data is useful in real-world HMA applications beyond Cambodia Landmine and Cluster Munition Monitor (2023); United Nations Development Programme in Cambodia (2021). While there is evidence that ontology-guided methods are effective (Cauter and Yakovets, 2024), they are limited to single-sentence inputs and overlook context-level reasoning. TextMine addresses this gap by enabling context-level reasoning, while aggregating a set of HMA-related ontologies. Finally, no framework is bias-aware, which might distort the evaluation results, a gap that TextMine also addresses.
To sum up, our contributions are: 1 Data: we introduce a curated dataset, curated ontology and annotated ground-truth data for humanitarian demining operations, systematically categorizing operational entities and relationships. 2 Evaluation: We evaluate extracted triples against our annotated dataset and introduce a bias-aware LLM-as-Judge framework for reference-free scoring. Experiments on closed and open LLMs show that position bias skews rankings. 3 Knowledge Extraction: We propose a prompt-based pipeline that combines layout-aware document chunking, ontology-guided extraction, and multi-perspective evaluation. To our knowledge, this is the first LLM application of knowledge extraction for HMA. 4 In-Context Learning Optimization: We found that prompts enriched with ontology-aligned examples improve triple extraction accuracy by up to 44.2%, reduce hallucinations by 22.5%, and enhance format conformance by up to 20.9% compared to baseline LLMs. These findings provide practical insights for prompt construction.
2 Benchmark Dataset Creation
We illustrate the process of creating the annotated humanitarian dataset in Fig. 2. We curated 120 online available technical reports: 60 reports from the Cambodian Mine Action Center (CMAC) and 60 from Geneva International Centre for Humanitarian Demining (GICHD) websites 1. The reports are all in the PDF format. The curation is based on relevance to the humanitarian demining, language (English), and recency. From the initial humanitarian dataset of 120 PDF documents 1, we filtered a dataset of five most recent mine action reports from CMAC of overall 233 pages (see Table 1) 2.
Table 1: Dataset Statistics for TextMine Corpus
Document Pages Chars Words Sent. Nums
Annual progress report 170 347,249 49,164 1,887 3,672
Mine clearance report 14 57,537 8,908 452 1,222
Integrated work plan 21 44,952 6,153 323 611
Cluster munition remnant report 9 37,846 5,865 327 893
Article 7 report 19 35,139 5,003 164 920
Total 233 522,723 75,093 3,153 7,318
Simultaneously, we selected domain-specific ontologies through working with domain experts, by first incorporating data models from the Information Management System for Mine Action (IMSMA) and then manually filtering them for high quality 3. We performed a survey of demining related ontologies together with domain experts and we incorporate six ontologies from Information Management System for Mine Action (IMSMA) Core [^2] https://www.gichd.org/our-response/information-management/imsma-core/. These ontologies are used for demining information system but not strictly for HMA, We then manually filter out concepts and relations unrelated to HMA. Furthermore, we add a more general humanitarian domain ontology from Empathi (Gaur et al., 2019) to make the overall HMA ontology more comprehensive. As a result, HMA ontology integrates seven ontologies (160 entity types, 86 relation types) covering diverse aspects of HMA (see details in Appendix B)..
The annotation prompt generation 4 first parses reports into text chunks, then systematically combines each chunk with ontology templates. Each template specifies the relation types for a specific domain (e.g., mine action events, land contamination, mine clearance). This generates 2,520 prompts (360 chunks $\times$ 7 templates), where each prompt contains both report context and ontology specifications for that domain. We randomly select 100 prompts 5 out of the 2,520, and apply them with the same prompts using GPT-4o and Llama3-70B 6 to initially annotate the data as the “LLM annotations 7.
For each prompt and data chunk pair, each LLM generates a set of knowledge triples as outputs. Then, a human annotator reviewed these outputs, filtering out incorrect knowledge triples and aggregating the valid ones in a new set. Incorrect knowledge triples can include triples following wrong format, mixed order of entities, hallucinated entities/relations (e.g., triples that are not included in the data chunk) and so on. In addition, there exist duplicate triples which are removed to create the final reference set. Those included 1,095 unique triples across the 100 prompts 8. The end result of the human annotation is a set of knowledge triples which are relevant and correct without hallucination or formatting problems, faithful to the given data chunk. This clean set comprises the “annotated test data” 9. The annotated test data can serve as the ground-truth for evaluations of performance of different techniques and LLMs.
Figure 2: Semi-automatic creation of the humanitarian mine action dataset. The dataset contains a large and diverse set of technical reports from the global mine action and a smaller LLM- and human-annotated portion for Cambodian mine action.
Finally, to assess the reliability of the annotations, a subject-matter expert—working in the demining innovation area—annotated 20 sets of data (corresponding to 286 knowledge triples) following the same methodology. Because we are interested in the overlap between the triples extracted by two independent annotators, we quantify the inter-annotator agreement with Jaccard Index, Dice Coefficient and Overlap Coefficient for these 20 sets. The two annotators have on average 0.89, 0.94, and 0.97 agreements respectively. These consistently high agreement scores indicate that the annotations in the full dataset are of similarly high quality.
@@ 3 TextMine Overview
3.1 Humanitarian Mine Action Task
Knowledge triple extraction from HMA reports can be formally defined as follows: given an ontology $\mathcal{O}=(\mathcal{E},\mathcal{R})$, where $\mathcal{E}$ is a set of entities and $\mathcal{R}$ is a set of relations, and a textual context $C$, the objective is to design a prompt $P(C,\mathcal{O})$ that guides an extractor model $M$ to extract triples $T=\{(s_{1},r_{1},o_{1}),\cdots,(s_{n},r_{n},o_{n})\mid s,o\in\mathcal{E},\,r\in\mathcal{R}\}$, where $n$ is the number of extracted triples, and subjects/objects ($s_{i}/o_{i}$), and relations ($r_{i}$) are extracted from the source text and mapped to $\mathcal{E}$ and $\mathcal{R}$ respectively. The extracted triples must remain consistent with both the ontology and the source text; i.e.  $T=M\bigl(P(C,\mathcal{O})\bigr)$.
Figure 3: TextMine Overview. Reports are preprocessed into paragraph chunks, used as test inputs during inference. Each chunk is combined with an instruction template and ontology, then passed through LLMs for triple extraction. We apply a multi-perspective evaluation using both reference-based and reference-free methods. Extracted triples are stored in a database, queried by developers and HMA domain experts.
3.2 Overall Pipeline
@@
Layout-Aware Document Chunking
@@
Ontology-Guided Knowledge Extraction
Given the text chunks as input, our goal is to extract $(s,r,o)$-triples. First, the extracted triples must be accurate w.r.t. the source text. An ideal extraction system would extract all triples (recall=1) precisely (precision=1). Second, extracted triples must conform to a specified ontology so that extracted demining knowledge can be stored and shared between organizations in a compatible way. We address this by combining a domain-specific HMA ontology with LLM in-context learning to extract triples from paragraphs.
@@
We design five prompting strategies for knowledge triple extraction: (1) Zero-shot (instruction and context only), (2) One-shot with Random Sentence (RS), (3) One-shot with Random Paragraph (RP), (4) One-shot with Ontology-Aligned Sentence (OS), and (5) One-shot with Ontology-Aligned Paragraph (OP). One example prompt is provided in Appendix.
Ontology-aligned demonstrations share the same ontology (entity types and relations) as the target context, while random demonstrations use unrelated ontology. Sentences are extracted from paragraphs using NLTK sentence tokenizer [^3] https://www.nltk.org/howto/stem.html. For OS and OP prompts, we design a retrieval algorithm that selects demonstrations from our annotated dataset (§ 2) by identifying the shortest context-answer pair matching the target ontology, for minimizing token costs while retaining high semantic similarity. To prevent data leakage and ensure a fair evaluation, we implement a second retrieval step if the initially-retrieved demonstration contains the same or part of target context as the test instance. In such cases, we select the next shortest matching example instead. This ensures that the retrieved demonstrations do not overlap with the evaluation context, preserving the integrity of the inference.
We design these prompt templates to test two hypotheses: (a) Ontology alignment enhances accuracy by priming the model with ontology-specific reasoning. Semantically aligned demonstrations help constrain the label space and improve precision (Min et al., 2022; Long et al., 2024). This is evident in our results, as the contrast between RS/RP and OS/OP confirms the benefit of ontology alignment. (b) Paragraph-level context improves extraction performance by providing richer demonstrations that reflect how entities and relations are introduced across sentences in real-world reports. However, our results do not support this hypothesis as comparisons between RS vs. RP and OS vs. OP show no consistent improvement from paragraph-level context.
@@ 3.3 Multi-Perspective Evaluation
HMA is a high-stakes decision-making domain, where incorrect triples—especially hallucinated ones—can misinform demining operations, leading to inefficiency. A comprehensive evaluation of triple extraction requires assessing accuracy, reliability, and structural validity. We evaluate models across three dimensions: (1) Triple Extraction Accuracy, (2) Hallucination Rate, and (3) Format Conformance.
@@
Triple extraction accuracy serves as the primary metric, as it directly measures how well models extract knowledge triples from text. However, accuracy alone does not fully capture model reliability. The hallucination rate evaluates faithfulness by detecting extraneous or fabricated information, while format conformance ensures that outputs adhere to a syntactically valid structure, enabling seamless integration into downstream applications. We employ N-gram matching-based metrics such as BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), and METEOR (Banerjee and Lavie, 2005) to assess the extracted triples. Additionally, we incorporate BERTScore (Zhang et al., 2019), which leverages word embeddings to capture semantic similarity beyond lexical overlap.
To enhance the accuracy of our metrics, we first apply stemming and lemmatization by NLTK to normalize morphological variations. We then compute the accuracy metrics against the manually annotated test set.
@@
Accuracy measures how well the extracted triples match reference triples, but it does not fully capture whether the generated content is grounded in the input. A model could produce plausible triples that are semantically similar to the original text—thus achieving a high accuracy score—yet still be incorrect; i.e. they do not appear in the input but seem reasonable. Hallucination is a prevalent issue (Ji et al., 2023; Huang et al., 2023; Xu et al., 2024b) and a critical aspect of our evaluation. To quantify this, for each extracted triple $t=(s,r,o)$ we first normalize $s$, $r$, and $o$, as well as the entire input context (tokenization, lemmatization, lowercasing, and punctuation removal). We flag $t$ as a hallucination if the normalized subject $s$ and object $o$ are not found as contiguous substrings in the normalized report text, or if the normalized relation $r$ is absent from the ontology set. This procedure ensures that even “plausible” but unsupported triples are detected and penalized, thereby maintaining faithfulness and trustworthiness.
@@
Format conformance metric assesses whether the generated triples adhere to the correct syntactic format of $r(s,o)$, where $r$ is the relation and $s$ and $o$ denote the subject and object, respectively. We consider a triple well-formatted if it follows this structure. We accommodate edge cases where the subject or object contains numerical values with commas, such as hasReliabilityInfo(2,500,011 square meters, landmine/ERW affected areas), or phrases in parentheses, such as hasAccidentOrganisationInfo(Quality of Life Survey (QLS), Department of Victim Assistance of CMAA). Format conformance ensures that extracted triples follow a structured format necessary for practical use. A model with high accuracy but poor format conformance may fail to produce usable outputs, limiting its applicability in real-world.
@@
where $S^{\prime}$ represents the normalized metric values within $[0,1]$, and the hallucination rate is inverted to penalize more hallucinations. $k=5$ is the number of metrics included in the score. Format conformance is excluded in the Combined Score, as our experimental results show consistently high format conformance across all extraction models and prompts, making it non-differentiating. Combined Score provides a holistic measure of extraction quality while mitigating scale differences among individual metrics.
@@
Evaluating generated texts is particularly challenging in domains like HMA, where annotated datasets are scarce. Demining reports are highly technical and domain-specific, requiring extracted triples to align with predefined ontologies of landmine types, clearance operations, and affected areas. Constructing a manually labeled test set is time-consuming and resource-intensive, limiting large-scale reference-based evaluation. Given these constraints, we explore an LLM-as-a-Judge approach as a potential reference-free evaluation framework for evaluating extracted triples. LLM-as-a-Judge offers a potential alternative to evaluation when ground-truth data is limited (Friel and Sanyal, 2023; Saad-Falcon et al., 2023; Es et al., 2023). The ultimate objective of our approach is to find an optimal judge LLM setting where the LLM consistently identifies the best candidate answer and provides a reasoned justification for its decision.
We try to find the optimal LLM Judge setting by conducting systematic ranking experiments and analyzing correlations between the LLMs judged ranking and reference-based rankings. For these ranking experiments, we design Judge Prompts that instruct LLMs on evaluation criteria. We use five models: Mistral-7B Jiang et al. (2023), Llama3-8B Grattafiori et al. (2024), Gemma2-9B Team et al. (2024), LLaMA3-70B and GPT-4o Hurst et al. (2024), as extraction models. We rank five responses from five models using GPT-4o, Llama3.1-70B, and Llama3.3-70B as our judge models. the correlation between different judge models and methods are included in Table 2. The judge prompts follow a fixed template with seven placeholders, where the ontology placeholder represents entity and relation types. Formally, let the input set for the LLM judge prompts be $\{O,C,R_{m_{1}},R_{m_{2}},R_{m_{3}},R_{m_{4}},R_{m_{5}}\}$, where $O$ is the ontology set, $C$ is the set of test contexts, and $R_{m_{1}},\cdots,R_{m_{5}}$ are the five sets of candidate answers from the five extractor models. The judge LLM produces a verdict (output as a ranking):
@@
Once the optimal judge LLM setting is determined, we adopt it as the reference-free evaluator to identify the best answer from each extraction, leveraging its reasoning process. Detailed prompt templates and an example of the reasoning process used for evaluation are provided in Appendix.